a rectangle is 1om longer than it is wide. its perimeter is 80m. find its length and width.
Answer by ReadingBoosters(3246)

10+w=l
80 = 2l+2w
Substitute the first equation into the second
80 = 2(10+w) + 2w
80 = 20 + 2w + 2w
80 - 20 = 20 - 20 + 4w
60 = 4w
60/4 = 4w/4
15 m = width
10+15=25 m = length